i have 71 coins in my wallet consisting of 10 peso, 5 peso coin, and 1 peso coins. if the 10 peso coins are 3 less than 3 times as many as 5 peso coins, and the 1 peso coins are 2 more than twice the 5 peso coins, how much money do i have in my wallet?
Answer by mananth(16946)

i have 71 coins in my wallet consisting of 10 peso, 5 peso coin, and 1 peso coins. if the 10 peso coins are 3 less than 3 times as many as 5 peso coins, and the 1 peso coins are 2 more than twice the 5 peso coins, how much money do i have in my wallet?
1 peso---------------------- 5 peso--------------------- 10 peso
2x+2---------------------------x--------------------------3x-3
2x+2 + x+ 3x-3 =71
6x=72
x= 12 5 peso coins
1 peso = 22
10 peso = 30
total money he has = 12*5 + 22*(2*12 +2) + 30*(3*12 -3)= 1622